In a advertisement, a pizza shop claims that its mean delivery time is less than 20...

  1. In a advertisement, a pizza shop claims that its mean delivery time is less than 20 minutes. of

    A random selection 50 delivery times has a sample mean of 22 minutes and a standard deviation

    of 2.1 minutes. Is there enough evidence to support the claim at = 0.005.

    Find:

  2. Question 5

    The null hypothesis ( H0)

    1. 22 min

    2. 22 min

    3. 20 min

    4. 20 min

  • Question 6

    The alternative hypothesis ( H0)

    1. 22 min

    2. 20 min

    3. 22 min

    4. 20 min

  • Question 7

    The critical value ( tC )

  • Question 8

    The test statistic ( tTEST

  • Question 9

    Your conclusion about the null hypothesis

    1. Fail to reject ( )

    2. Reject ( )

  • Question 10

    Your conclusion about the manufacturer's claim.

    1. At 0.5% level of significance, There is not enough evidence to conclude that the mean delivery time is less than 20 minutes.

    2. At 5% level of significance, there is enough evidence to conclude that the mean delivery time is less than 20 minutes.

    3. At 0.5% level of significance, there is not enough evidence to conclude that the mean delivery time is more than 20 minutes.

    4. At 05% level of significance, there is enough evidence to conclude that the mean delivery time is more than 20 minutes.
Answer #1

given data are:-

sample size(n) = 50

sample mean() = 22

sample standard deviation(s) = 2.1

5).The null hypothesis:-

6).The alternative hypothesis:-

7). df = (n-1) = (50-1) = 49

The critical value :-

[ using t distribution table, for f = 49,alpha= 0.005 , left tailed test]

8).test statistic be:-

9).conclusion:-

so, we fail to reject the null hypothesis.

10).conclusion about the manufacturer's claim is :-

At 0.5% level of significance, There is not enough evidence to conclude that the mean delivery time is less than 20 minute

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible...